OpenStudy (anonymous):

5/6x + 1/2 =1/4 (x-4)

OpenStudy (anonymous):

hi , how are you first??

OpenStudy (anonymous):

great and you???

OpenStudy (anonymous):

good

OpenStudy (anonymous):

can you give me steps to see how you solved this first?

OpenStudy (anonymous):

I need to learn I have a midterm coming up quick and I have to have this down

OpenStudy (anonymous):

5\6x+1\2=1\4x-1 5\6x-1\4x=-3\2 20x-6x\24x=-3\2 14x\24x=-3\2 7x\12x=-3\2 -14x-36x=0

OpenStudy (anonymous):

iam busy right now brb

OpenStudy (anonymous):

(5/6)x + 1/2 = 1/4 (x-4) multiply 1/4 by (x-4), you get: (5/6)x + 1/2 = (1/4)x - 1 put x's on one side and numbers on the other side: (5/6)x - (1/4)x = (-1/2) - 1 subtract (5/6) and (1/4) (7/12) x = -3/2 multiply both sides by 12/7 x = -18/7 done